# Mass-loaded magnetic explosions in the context of Magnetar Giant Flares and Fast Radio Bursts

### v1

@@ 3 Solutions
By taking the limit of $v\ll 1$ we find that there are two main classes of solutions of equation (22). A class where $g\propto v^{-1}$, leading to solutions that reduce to a magnetic dipole near the origin, and a class where $g\propto v^{2}$ where the field near the centre is approximately uniform. While the former case was explored in detail in Gourgouliatos & Vlahakis (2010), the latter has not been studied. In this work we will focus on the latter class of solutions that are not singular at the origin by exploring appropriate combinations of the main parameters $c_{0}$ and $c_{1}$.
The equation does not permit general analytical solutions and we shall proceed with a numerical solution. Given the singularities at $v=0$ and $v=1$, we integrate within an interval $v\in[v_{min},v_{max}]$, where $v_{min}=10^{-3}$ and $v_{max}<1$. We set $g(v_{min})^{\prime}=0$, and $g(v_{min})=g_{0}$, we choose $g_{0}$, so that the maximum value of $g$ is set to unity. This is straightforward for the homogeneous version of the equation with $c_{1}=0$, as a simple division with the maximum value of a trial solution with arbitrary $g_{0}$ leads to a normalised solution. However, in the non-homogeneous case ($c_{1}\neq 0$) it requires an iterative processes as simple division of the function does not lead to a solution. Thus, we repeat the process by dividing $g_{0}$ by the maximum value of the solution and integrating, until convergence. We use a tolerance limit of $10^{-4}$ for the normalisation and we typically find that the solution converges to a normalised form after less than $100$ iterations. For the integration of the ordinary differential equation we used the Runge-Kutta-Fehlberg method, applying the built-in routines of numpy and scipy in python (Harris et al., 2020; Virtanen et al., 2020).
@@
The results of the integration for $c_{1}=0$ and various values of $c_{0}$, ranging from $1$ to $100$ are shown in Figure 1, which are fully consistent with the analytical solutions reported in Barkov et al. (2022), with the note that the value of $\alpha$ there corresponds to $\sqrt{c_{0}}$ in this work. It can be seen clearly that by increasing the value of $c_{0}$ the maximum velocity attainted by the expanding field becomes smaller. This velocity corresponds to the first root, beyond the origin $v=0$, as $g(0)=0$, of the function $g(v_{0})=0$. Thus, by increasing the ratio of the poloidal current over the poloidal field, or equivalently the ratio of the toroidal field over the poloidal field, the maximum expansion velocity is smaller. We note that beyond the $v_{0}$ the solutions becomes oscillatory, with the flux taking negative and positive values. However, this part of the solution is completely disconnected from the inner part, both in terms of the poloidal and the toroidal fields, thus no field lines from the central part are connected to the outer ones. Thus we halt the integration at $v_{0}$. These solutions are denoted as Z-type solutions, as the pressure and density are zero.
@@
Next, we have explored solutions with non-zero values of $c_{1}$. Due to the large number of combinations, we have decided to keep the toroidal field parameter constant ($c_{0}=1$) and then explore positive values of $c_{1}$, ranging from $0.1$ to $100$, with the results shown in Figure 2. The inclusion of pressure and density leads to a structure that is less extended in the velocity space, reaching a smaller maximum velocity compared to the $c_{1}=0$ case. We have denoted these solutions as P-type solutions, as the pressure and density increase with increasing magnetic flux function.
@@
Finally, it is possible to explore negative values of $c_{1}$. These cases correspond to plasmoids where the pressure decreases as the magnetic flux function increases. Despite this negative relation, the freedom of an additive constant $\tilde{p}_{0}$ in equation (21) prevents from negative values of the pressure and the density. These solutions have the remarkable property, that for an appropriate combination of $c_{0}$ and $c_{1}$, not only the flux function reaches 0 at $v_{0}$ but also its derivative becomes zero there thus: $g^{\prime}(v_{0})=0$, as shown in Figure 3. This has the important consequence that all three components the magnetic field are zero on the surface of the sphere. We note that if the values of $c_{1}$ are smaller algebraically than the ones corresponding to the zero derivative on the boundary, the solutions become unphysical with the field lines not closing but increasing infinitely. The resulting structures correspond to underdensities, where the drop of thermal plasma pressure and mass density within the sphere is compensated by the magnetic and electric field pressure and tension. We have denoted these as N-type solutions as in these solutions the density and pressure decrease with increasing flux function.
@@
The parameters provided in table LABEL:tab:1 were employed for the three families of solutions: $c_{1}=0$ Z-models, where there is no mass or thermal pressure, $c_{1}>0$ P-models, where there is mass and pressure varying with the same sign of the flux function, and $c_{1}<0$ N-models, where, in addition to the negative sign of $c_{1}$, the derivative of $g$ at $v_{0}$ vanishes.
@@
Table 1: Solution properties: The columns represent the solution number, parameters $c_{0}$, $c_{1}$, the root of the solution $v_{0}$ and the corresponding Lorentz number $\Gamma_{0}$, the magnetic field energy $E_{mag}$, the electric field energy $E_{el}$, the work of the pressure $W$ and the mass $M$ enclosed in the solution. The units of the various energy terms ($E_{mag}$, $E_{el}$ and $W$) are in units of $10^{45}$ erg and the mass in units of $1.8\times 10^{26}$g, under the assumption of a sphere of were $v=1$, corresponds to a radius of 11km the magnetic flux $g$ is normalised to $10^{25}$G cm2.
@@
Figure 1: The Z-family solutions for $g(v)$, for various values of $c_{0}$ while $c_{1}=0$.
Figure 2: Characteristic P-family solutions of $g(v)$ various values of $c_{1}$, while keeping $c_{0}=1$.
Figure 3: The N-family of solutions showing $g(v)$, for various values of $c_{0}$ and negative $c_{1}$, so that the derivative is zero at the maximum velocity.
@@ 4.1 General properties of the solutions
Below we present a description of the main characteristics of the solutions depending on the values of the parameters used.
@@
For all solutions, smaller values of $c_{0}$ lead to $v_{0}$ that approaches unity, thus the system becomes more relativistic. This implies that the more twisted the solution is, or equivalently if the poloidal current is higher, the plasmoid is contained within a smaller sphere. This property is related to the fact that by assuming a constant $c_{0}$ a particular length-scale is introduced. In this case the units of $\sqrt{c_{0}}$ are inversely proportional to length. Consequently, by increasing $c_{0}$ this length-scale becomes smaller. This is consistent with the well-known non-relativistic, linear force-free axisymmetric solutions in spherical geometry (Chandrasekhar & Kendall, 1957) whose radius shifts to smaller values for larger toroidal-to-poloidal magnetic field ratio. This property is well-preserved and reflected on its relativistic analogue.
@@
Regarding the surface terms, in Z and P solutions, there exists a non-vanishing meridional magnetic field on the surface $v=v_{0}$. Because of this magnetic field discontinuity an azimuthal electric current sheet flows on the surface of the plasmoid. Moreover, there also exists a volume electric charge, due to the divergence of the meridional component of the electric field. Following the convention of a positive value for the flux function, the charge is positive in the northern hemisphere and negative in the southern one. While this structure is in equilibrium at its interior, on its surface the magnetic pressure is anisotropic, stronger at the equator and weaker at the poles. Therefore, if it is embedded within a medium of uniform pressure, it is expected to expand on the equatorial direction, while the presence of the surface current may trigger resistive instabilities and lead to reconnection on a layer that is propagating relativistically.
@@
Considering the solutions with $c_{1}>0$, we find that higher values of $c_{1}$ lead to plasmoids with smaller $v_{0}$. This is due to the inclusion of extra pressure and density that prevent the system from extending to a $v_{0}$ as high as in the purely electromagnetic case ($c_{1}=0$). The overall structure has similar characteristics to the previous solution, regarding the surface azimuthal current sheet and pressure profile and it be prone to similar instabilities.
@@
Solutions of the N family, with $c_{1}<0$ lead to different qualitative properties. In general choosing a negative $c_{1}$ leads to $v_{0}$ closer to unity compared to the solutions with zero or positive $c_{1}$ and the same $c_{0}$. Below a critical value of $c_{1}$, the solution has no roots, thus becoming unphysical. In the N-class of solutions we have determined these critical values for various choices of $c_{0}$. The solutions corresponding to these critical values have the remarkable property that at $v_{0}$, not only the function $g$ is zero but also its derivative vanishes. This implies that all three components of the magnetic field, and both components of the electric field are zero on the spherical surface defined by $v=v_{0}$. Because of this, there is no anisotropy nor a surface current due to the meridional component. Thus, if spherical plasmoids obeying these solutions are embedded within a medium of uniform pressure the system can maintain its equilibrium, as opposed to structures corresponding to the other families of solutions. Moreover, the lack of surface current sheets makes less likely the appearance of current driven instabilities.
@@ 4.2 Physical Quantities
Based on these solutions it is possible to calculate the corresponding physical quantities arising from these calculations. In what follows we focus on the electromagnetic field energy, the mass of the sphere, the work of the thermal pressure and the temperature of the plasmoid. A set of characteristic configurations and the corresponding physical properties are shown in Figures 4 and 5.
@@
Figure 4: Contour plots of the solutions with $c_{0}=1$ and $c_{1}=1$ (top row); and $c_{0}=1$ and $c_{1}=10$ (bottom row). The black contours correspond to the poloidal magnetic field lines, while in colour is plotted: the $B_{\phi}$ component of the magnetic field (panels a, e), the rest mass density $\rho$ (panels b and f), the temperature $T$ (panels c and g); and the electric charge density $j^{0}$ (panels d and h). The solution remains invariant in the $v$-space, here we present a snapshot at some time $t$, so that $r=1$ corresponds to $v=1$.
@@
Figure 5: Contour plots of the solutions with $c_{0}=10$ and $c_{1}=-1.448$ (top row); and $c_{0}=100$ and $c_{1}=-211.3$ (bottom row). The black contours correspond to the poloidal magnetic field lines, while in colour is plotted: the $B_{\phi}$ component of the magnetic field (panels a, e), the rest mass density $\rho$ (panels b and f), the temperature $T$ (panels c and g); and the electric charge density $j^{0}$ (panels d and h). As in the plots of Figure 5, the snapshot is taken at some time $t$, so that $r=1$ corresponds to $v=1$.
@@ 4.2.1 Electromagnetic Field
The electric and magnetic field are given from equations (5) and (12) with appropriate substitution of the solutions and the velocity profile. The electromagnetic field is contained within a sphere in the $v$-space. The poloidal field is approximately uniform close to the axis, and closes within a expanding sphere of radius $r_{0}=cv_{0}\penalty 10000\ t$. The $B_{\phi}$ field peaks at the equatorial plane. However this peak does not coincide with the maximum of the flux function $\Psi$, which corresponds to the null point of the closed poloidal loops, Figures 4 and 5 panels a and e. This is due to the fact that the $B_{\phi}$ is proportional to $\Psi$, but is also multiplied by a factor $\frac{v\Gamma^{2}}{r^{2}}$, which for the highly relativistic case dominates and pushes the maximum to more distant radii. Therefore, in the highly relativistic case, a thin layer containing a toroidal field expanding relativistically is expected, while in the non-relativistic case $v_{0}\ll 1$ the toroidal field is contained closer to the centre.
We can then calculate the energy of the electric and the magnetic field:
@@
From the above expression it clear that the magnetic field energy stored in the spherical configuration is inversely proportional to time, decreasing as the radius expands.
@@
As the electric field is the $\theta$ and $\phi$ components of the magnetic field multiplied by the velocity, it follows a similar behaviour, with electric field energy being inversely proportional to time. Thus, as the system expands it decreases. Electric and magnetic field energy contained within a sphere as a function of $c_{0}$ is shown in figure 6, where the solution for the Z ($c_{1}=0$) and N-models ($c_{1}<0)$ is shown. We find that for smaller and larger values of $c_{0}$ the magnetic field energy is higher and has a minimum for intermediate $c_{0}$. We remark that in the solutions we have normalised the maximum value of $g$ and consequently $\Psi$ to unity. Because of this, while the magnetic flux is the same in all solutions its energy content is related to how it is distributed in the sphere. In the case of a smaller sphere and large $c_{0}$, all components of the magnetic field are large since they scale inversely with $r$. For spheres where $v_{0}$ approaches unity, while the radius is higher, $g$ develops a sharp derivative at $v_{0}$, leading to a strong $B_{\theta}$ component and energy content. On the contrary, the electric field energy has a monotonic dependence $c_{0}$ as its intensity is proportional to $v$, therefore for smaller spheres, in velocity space, its overall contribution is smaller, even though the magnetic field energy is higher.
Figure 6: The electric and magnetic energy for the the Z and N solutions as functions of $c_{0}$.

### v2

@@ 3 Solutions
The behaviour of the solution near the origin can be inspected for $v\ll 1$. The leading terms of the equation are the following:
$v^{2}g^{\prime\prime}-2g=0\,.$ (23)
This admits solutions in following form:
$g(v)=A_{1}v^{2}+A_{2}v^{-1}\,$, (24)
@@
where $A_{1}$ and $A_{2}$ are constants. Thus, there are two main classes of solutions of Eq. (22). A class where $g\propto v^{-1}$, leading to solutions that reduce to a magnetic dipole near the origin, and a class where $g\propto v^{2}$, where the field near the centre is approximately uniform. While the former case was explored in detail in Gourgouliatos and Vlahakis (2010), the latter has not been studied. In this work, we focus on the latter class of solutions, which are not singular at the origin, by exploring appropriate combinations of the main parameters $c_{0}$ and $c_{1}$.
3.1 Non-relativistic limit
In the non-relativistic limit, while keeping the terms proportional to $c_{0}$ and $c_{1}$, Eq. (22) reduces to
$v^{2}g^{\prime\prime}-2g+c_{0}v^{2}g+c_{1}v^{4}=0\,.$ (25)
The solution of the above equation is
$g(v)$ $=$ $B_{1}\left(\frac{\sin\left(\sqrt{c_{0}}v\right)}{\sqrt{c_{0}}v}-\cos\left(\sqrt{c_{0}}v\right)\right)$ (26)
$+$ $B_{2}\left(\frac{\cos\left(\sqrt{c_{0}}v\right)}{\sqrt{c_{0}}v}+\sin\left(\sqrt{c_{0}}v\right)\right)-\frac{c_{1}}{c_{0}}v^{2}\,$,
where $B_{1}$ and $B_{2}$ are constants. The term multiplied with $B_{1}$ is zero at the origin, whereas the term multiplied with $B_{2}$ is singular. Therefore, we kept the $B_{1}$ terms and set $B_{2}=0$. This solution is mathematically identical to the solutions found in Gourgouliatos et al. (2010, Eq. 6), where the role of $v$ is taken by $r$. In the present case, the sphere instead of being static, it expands uniformly.
3.2 Relativistic force-free limit
@@
In the absence of pressure and inertia, we have $c_{1}=0$. The results of the integration for $c_{1}=0$ and various values of $c_{0}$, ranging from $1$ to $100$, are shown in Fig. 1. They are fully consistent with the analytical solutions reported by Barkov et al. (2022), with the note that the value of $\alpha$ there corresponds to $\sqrt{c_{0}}$ in this work. Clearly, when the value of $c_{0}$ increases, the maximum velocity attained by the expanding field becomes lower. This velocity corresponds to the first root, beyond the origin $v=0$, as $g(0)=0$, of the function $g(v_{0})=0$. Thus, by increasing the ratio of the poloidal current over the poloidal field, or equivalently, the ratio of the toroidal field over the poloidal field, the maximum expansion velocity is lower. We note that beyond $v_{0}$, the solutions become oscillatory, with the flux taking negative and positive values. However, this part of the solution is completely disconnected from the inner part in terms of the poloidal and the toroidal fields, and thus, no field lines from the central part are connected to the outer ones. Thus, we halted the integration at $v_{0}$. These solutions are denoted as Z-type solutions, as the pressure and density are zero.
3.3 General solution
When all terms are included so that $c_{0}\neq 0$ and $c_{1}\neq 0$ and the system is relativistic, the equation does not permit general analytical solutions, and we proceed with a numerical solution. Given the singularities at $v=0$ and $v=1$, we integrated within an interval $v\in[v_{min},v_{max}]$, where $v_{min}=10^{-3}$ and $v_{max}<1$. We set $g(v_{min})^{\prime}=0$ and $g(v_{min})=g_{0}$, and we chose $g_{0}$ so that the maximum value of $g$ was set to unity. This is straightforward for the homogeneous version of the equation with $c_{1}=0$, as a simple division with the maximum value of a trial solution with arbitrary $g_{0}$ leads to a normalised solution. However, in the non-homogeneous case ($c_{1}\neq 0$), it requires an iterative processes as simple division of the function does not lead to a solution. Thus, we repeated the process by dividing $g_{0}$ by the maximum value of the solution and integrated until convergence. We used a tolerance limit of $10^{-4}$ for the normalisation, and we typically found that the solution converged to a normalised form after fewer than $100$ iterations. For the integration of the ordinary differential equation, we used the Runge-Kutta-Fehlberg method, applying the built-in routines of numpy and scipy in Python (Harris et al., 2020; Virtanen et al., 2020).
@@ 3.3 General solution
Next, we explored solutions with non-zero values of $c_{1}$. Because of the very many possible combinations, we decided to keep the toroidal field parameter constant ($c_{0}=1$) and then explore positive values of $c_{1}$ ranging from $0.1$ to $100$. The results are shown in Fig. 2. The inclusion of pressure and density leads to a structure that is less extended in the velocity space and reaches a lower maximum velocity than in the $c_{1}=0$ case. We denoted these solutions P-type solutions, as the pressure and density increase with increasing magnetic flux function.
@@
Finally, it is possible to explore negative values of $c_{1}$. These cases correspond to plasmoids where the pressure decreases as the magnetic flux function increases. Despite this negative relation, the freedom of an additive constant $\tilde{p}_{0}$ in Eq. (21) prevents negative values of the pressure and density. These solutions have the remarkable property that for an appropriate combination of $c_{0}$ and $c_{1}$, not only the flux function reaches 0 at $v_{0}$, but its derivative also becomes zero there, and thus, $g^{\prime}(v_{0})=0$, as shown in Fig. 3. This has the important consequence that all three components of the magnetic field are zero on the surface of the sphere. We note that when the values of $c_{1}$ are smaller algebraically than those corresponding to the zero derivative on the boundary, the solutions become unphysical, with the field lines not closing, but increasing infinitely. The resulting structures correspond to underdensities, where the drop of thermal plasma pressure and mass density within the sphere is compensated for by the magnetic and electric field pressure and tension. We denoted these as N-type solutions as in these solutions the density and pressure decrease with increasing flux function.
@@
The parameters provided in Table LABEL:tab:1 were employed for the three families of solutions: $c_{1}=0$ Z-models, without mass or thermal pressure, $c_{1}>0$ P-models, with varying mass and pressure with the same sign as the flux function, and $c_{1}<0$ N-models, where, in addition to the negative sign of $c_{1}$, the derivative of $g$ at $v_{0}$ vanishes.
@@
Table 1: Solution properties.  [^1] The columns represent the solution number, parameters $c_{0}$ and $c_{1}$, the root of the solution $v_{0}$, and the corresponding Lorentz number $\Gamma_{0}$, the magnetic field energy $E_{mag}$, the electric field energy $E_{el}$, the work of the pressure $W$, and the mass $M$ enclosed in the solution. The units of the various energy terms ($E_{mag}$, $E_{el}$, and $W$) are in units of $10^{45}$ erg and the mass is in units of $1.8\times 10^{26}$g, under the assumption of a sphere where $v=1$ corresponds to a radius of 11km, and the magnetic flux $g$ is normalised to $10^{25}$G cm2.
@@
Figure 1: Z-family solutions for $g(v)$ for various values of $c_{0}$ when $c_{1}=0$.
Figure 2: Characteristic P-family solutions of $g(v)$ for various values of $c_{1}$ while keeping $c_{0}=1$.
Figure 3: N-family of solutions showing $g(v)$ for various values of $c_{0}$ and negative $c_{1}$, so that the derivative is zero at the maximum velocity.
@@ 4.1 General properties of the solutions
We describe the main characteristics of the solutions depending on the values of the parameters below. For all solutions, lower values of $c_{0}$ lead to a $v_{0}$ that approaches unity, and thus, the system becomes more relativistic. This implies that the more twisted the solution, or equivalently, the higher the poloidal current, the smaller the sphere containing the plasmoid. This property is related to the fact that by assuming a constant $c_{0}$, a particular length-scale is introduced. In this case, the units of $\sqrt{c_{0}}$ are inversely proportional to the length. Consequently, by increasing $c_{0}$, this length scale becomes smaller. This is consistent with the well-known non-relativistic linear force-free axisymmetric solutions in spherical geometry (Chandrasekhar and Kendall, 1957), whose radius shifts to lower values for a higher toroidal-to-poloidal magnetic field ratio. This property is well preserved and reflected on its relativistic analogue.
@@
In the Z and P solutions, a non-vanishing meridional magnetic field exists on the surface $v=v_{0}$. Because of this magnetic field discontinuity, an azimuthal electric current sheet flows on the surface of the plasmoid. Moreover, there also exists a volume electric charge because the meridional component of the electric field diverges. Following the convention of a positive value for the flux function, the charge is positive in the northern and negative in the southern hemisphere. While this structure is in equilibrium at its interior, the magnetic pressure is anisotropic on its surface, stronger at the equator, and weaker at the poles. Therefore, when it is embedded within a medium of uniform pressure, it is expected to expand in the equatorial direction, while the surface current might trigger resistive instabilities and lead to reconnection in a layer that propagates relativistically.
@@
Considering the solutions with $c_{1}>0$, we find that higher values of $c_{1}$ lead to plasmoids with smaller $v_{0}$. This is due to the inclusion of additional pressure and density that prevent the system from extending to a $v_{0}$ as high as in the purely electromagnetic case ($c_{1}=0$). The characteristics of the overall structure are similar to those of the previous solution regarding the surface azimuthal current sheet and pressure profile, and the structure is prone to similar instabilities.
@@
Solutions of the N family with $c_{1}<0$ lead to different qualitative properties. In general, choosing a negative $c_{1}$ leads to $v_{0}$ closer to unity than in solutions with zero or positive $c_{1}$ and the same $c_{0}$. Below a critical value of $c_{1}$, the solution has no roots and thus becomes unphysical. In the N-class of solutions, we determined these critical values for various choices of $c_{0}$. The solutions corresponding to these critical values have the remarkable property that at $v_{0}$, not only is the function $g$ zero, but its derivative also vanishes. This implies that all three components of the magnetic field and the two components of the electric field are zero on the spherical surface defined by $v=v_{0}$. Because of this, there is no anisotropy or a surface current due to the meridional component. Thus, when spherical plasmoids obeying these solutions are embedded within a medium of uniform pressure, the system can maintain its equilibrium, as opposed to structures corresponding to the other families of solutions. Moreover, the lack of surface current sheets makes the appearance of current-driven instabilities less likely.
@@ 4.2 Physical quantities
Based on these solutions, it is possible to calculate the corresponding physical quantities arising from these calculations. In what follows, we focus on the electromagnetic field energy, the mass of the sphere, the work of the thermal pressure, and the temperature of the plasmoid. A set of characteristic configurations and the corresponding physical properties are shown in Figs. 4 and 5.
@@
Figure 4: Contour plots of the solutions with $c_{0}=1$ and $c_{1}=1$ (top row) and $c_{0}=1$ and $c_{1}=10$ (bottom row). The black contours correspond to the poloidal magnetic field lines, and we plot in colour the $B_{\phi}$ component of the magnetic field (panels a and e), the rest-mass density $\rho$ (panels b and f), the temperature $T$ (panels c and g), and the electric charge density $j^{0}$ (panels d and h). The solution remains invariant in the $v$-space. We present a snapshot at some time $t$, so that $r=1$ corresponds to $v=1$.
@@
Figure 5: Contour plots of the solutions with $c_{0}=10$ and $c_{1}=-1.448$ (top row) and $c_{0}=100$ and $c_{1}=-211.3$ (bottom row). The black contours correspond to the poloidal magnetic field lines, and we plot in colour the $B_{\phi}$ component of the magnetic field (panels a and e), the rest-mass density $\rho$ (panels b and f), the temperature $T$ (panels c and g), and the electric charge density $j^{0}$ (panels d and h). As in the plots of Fig. 5, the snapshot is taken at some time $t$, so that $r=1$ corresponds to $v=1$.
@@ 4.2.1 Electromagnetic field
The electric and magnetic fields are given from Eqs. (5) and (12) with appropriate substitution of the solutions and the velocity profile. The electromagnetic field is contained within a sphere in the $v$-space. The poloidal field is approximately uniform close to the axis, and it closes within an expanding sphere of radius $R=cv_{0}\penalty 10000\ t$. The $B_{\phi}$ field peaks at the equatorial plane. However, this peak does not coincide with the maximum of the flux function $\Psi$, which corresponds to the null point of the closed poloidal loops (Figs. 4 and 5 panels a and e) because $B_{\phi}$ is proportional to $\Psi$, but is also multiplied by a factor $\frac{v\Gamma^{2}}{r^{2}}$, which for the highly relativistic case dominates and pushes the maximum to more distant radii. Therefore, in the highly relativistic case, a thin layer containing a toroidal field expanding relativistically is expected, while in the non-relativistic case $v_{0}\ll 1$, the toroidal field is contained closer to the centre.
We then calculated the energy of the electric and the magnetic field,
@@
It follows from Eq. (27) that the magnetic field energy scales inversely with time. This behaviour can be understood from simple scaling arguments: under self-similar expansion, the magnetic flux is conserved, implying that the characteristic magnetic field strength scales as $B\propto R^{-2}$. Since the magnetic energy density scales as $B^{2}\propto R^{-4}$, integration over the expanding spherical volume ($\propto R^{3}$) yields a total magnetic energy $E_{\rm mag}\propto R^{-1}$. Because the outer radius of the configuration evolves as $R=cv_{0}t$, this directly leads to $E_{\rm mag}\propto t^{-1}$.
@@
As the electric field is the $\theta$ and $\phi$ components of the magnetic field multiplied by the velocity, it follows a similar behaviour, with the electric field energy being inversely proportional to time. Thus, as the system expands, it decreases. We show the electric and magnetic field energy contained within a sphere as a function of $c_{0}$ in Fig. 6, where the solution for the Z ($c_{1}=0$) and N-models ($c_{1}<0)$ is shown. For lower and higher values of $c_{0}$, the magnetic field energy is higher and has a minimum for intermediate $c_{0}$. We remark that we normalised the maximum value of $g$ and consequently $\Psi$ to unity in the solutions. Because of this, while the magnetic flux is the same in all solutions, its energy content is related to its distribution in the sphere. In the case of a smaller sphere and large $c_{0}$, all components of the magnetic field are large since they scale inversely with $r$. For spheres where $v_{0}$ approaches unity, while the radius is higher, $g$ develops a sharp derivative at $v_{0}$, leading to a strong $B_{\theta}$ component and energy content. In contrast, the electric field energy has a monotonic dependence on $c_{0}$ as its intensity is proportional to $v$, and therefore, its overall contribution is smaller for smaller spheres in velocity space, even though the magnetic field energy is higher.
Figure 6: Electric and magnetic energy for the Z and N solutions as functions of $c_{0}$.